# Deep Class-Incremental Learning: A Survey

### v1

@@ 4 Experimental Evaluation
Sections 4.2 and 4.3 answer the first question, and Section 4.5 answers the second question. We provide the holistic performance measures in Section 4.6 to answer the third question. We also analyze the influence of incremental learning by visualizing the confusion matrix in Section 4.4 and summarize the comparison results in Section 4.7. Finally, we report more experimental results, implementation details, and visualizations of fully connected layers in the supplementary material.
@@
Figure 6: Incremental accuracy of different methods on CIFAR100. Legends are shown in (d).
@@ 4.1.1 Benchmark Datasets
iCaRL [32] firstly formulates the comparison protocol of class-incremental learning, which was widely followed and compared in other works. It suggests using CIFAR100 [31] and ImageNet100/1000 [190] for evaluation. They are listed as:
  • CIFAR100: There are 100 classes with 60,000 images, in which 50,000 are training instances and 10,000 are testing ones, with 100 images per class. Each image is represented by $\times32$ pixels.
  • ImageNet1000: is a large-scale dataset with 1,000 classes, with about 1.28 million images for training and 50,000 for validation. The spatial size of images varies.
  • ImageNet100: is the subset of ImageNet1000 containing 100 classes [32]. These classes are selected from the first 100 classes after a random shuffle.
There are other works [77, 75, 116, 92, 104, 139, 82] utilizing MNIST [191], Core50 [192], Google Landmark-v2 [193], CUB200-2011 [194], TinyImageNet [195], and miniImageNet [190] for model evaluation, while the aforementioned datasets are the most widely adopted in the current CIL community. Hence, we choose CIFAR100 and ImageNet100/1000 for model evaluation.
Table II: Average and last accuracy performance comparison on CIFAR100. ‘#P’ represents the number of parameters (million). The full table is reported in the supplementary.
@@
(a) ImageNet100 Base0 Inc5
(b) ImageNet100 Base0 Inc10
(c) ImageNet100 Base50 Inc10
(d) ImageNet100 Base50 Inc25
Figure 7: Incremental accuracy of different methods on ImageNet100. Legends are shown in (d).
Dataset Split: Following the protocol defined in [32], all classes are first shuffled by Numpy random seed $1993$. Afterward, there are two different ways to split the classes into incremental stages:
  • Train from scratch (TFS): splits all the classes equally into each incremental stage. For example, if there are $B$ stages and $C$ classes in total, each incremental task contains C/B{C}/{B} classes for training.
  • Train from half (TFH): splits half of the total classes as the first incremental task and equally assigns the rest classes into the following stages. Specifically, it assigns $C/2$ classes to the first task and C/2(B1)C/2(B-1) classes to the rest tasks.
Both of these settings are widely adopted in the current CIL community [93, 87]. Hence, to unify these settings, we use ‘Base-$m$, Inc-$n$’ to denote these settings where $m$ stands for the number of classes in the first stage, and $n$ stands for the number of classes in each incremental task. $m=0$ stands for the ‘Train from scratch’ protocol. We use the same training splits for every compared method for a fair comparison. The testing set is the same as the original one for holistic evaluation.
@@ 4.1.2 Evaluation Metrics
There are several metrics to evaluate the CIL model. We denote the Top-1 accuracy after the $b$-th task as $\mathcal{A}_{b}$, and higher $\mathcal{A}_{b}$ indicates a better prediction accuracy. Since the CIL model is continually updated, the accuracy often decays with more tasks incorporated. Hence, the accuracy after the last stage ($\mathcal{A}_{B}$) is a proper metric to measure the overall accuracy among all classes.
@@
On the other hand, only comparing the final accuracy ignores the performance evolution along the learning trajectory. Hence, another metric denoted as ‘average accuracy’ considers the performance of all incremental stages: $\bar{\mathcal{A}}=\frac{1}{B}\sum_{b=1}^{B}\mathcal{A}_{b}$. A higher average accuracy denotes a stronger performance along the incremental stages. Apart from Top-1 accuracy, some works address Top-5 accuracy for the ImageNet dataset. In the supplementary, we also report the Top-5 accuracy of these comparisons.
(a) ImageNet1000 Base0 Inc100
(b) ImageNet1000 Base500 Inc100
Figure 8: Incremental accuracy of different methods on ImageNet1000. Legends are shown in (b).
(a) Finetune
(b) EWC
(c) LwF
(d) Replay
(e) iCaRL
(f) BiC
(g) DER
(h) FOSTER
(i) MEMO
(j) Dytox
Figure 9: Confusion matrix of different methods on CIFAR100 Base0 Inc20 setting after the last incremental stage. The dotted red line splits the last 20 classes from the 80 classes learned in former tasks.
4.1.3 Implementation Details
Selected methods: In the comparison, we aim to contain all kinds of methods listed in the taxonomy of Table I. We systematically choose 16 methods, including: Replay [196], RMM [197] (data replay), GEM [59] (data regularization), EWC [77] (parameter regularization), FOSTER [21], MEMO [22], DER [20], DyTox [23], L2P [72] (dynamic networks), LwF [85], iCaRL [32], PODNET [98], Coil [92] (knowledge distillation), WA [115], BiC [87] (model rectify). We also report the baseline method ‘finetune,’ which updates the model with Eq. 4.
Among these methods, RMM is a specific technique to efficiently organize the memory budget, which can be orthogonally combined with other methods, and we combine it with FOSTER, denoted as FOSTER+RMM. L2P requires pre-trained ViT as the backbone model, while others are trained from scratch.
Training details: We implement the above methods with PyTorch [198] and PyCIL [199]. Specifically, we use the same network backbone for all CNN-based compared methods, i.e., ResNet32 [200] for CIFAR100 and ResNet18 for ImageNet. We use SGD with an initial learning rate of 0.1 and momentum of 0.9. The training epoch is set to 170 for all datasets with a batch size of 128. The learning rate suffers a decay of $0.1$ at 80 and 120 epochs. For ViT-based methods like DyTox and L2P, we follow the original implementation and use ConViT [201] for DyTox and pre-trained ViT-B/16 [24] for L2P. The optimization parameters of them are set according to the original paper since ViT has a different optimization preference to CNN. We follow the original paper to set the algorithm-specific parameters, e.g., splitting 10% exemplars from the exemplar set as validation for BiC, setting the temperature $\tau$ to 5 and using a 10 epochs warm-up for DER, using $\ell_{2}$ norm to normalize the fully-connected layers in WA.
Research has shown that a good starting point (i.e., the first stage accuracy $\mathcal{A}_{1}$) implies better transferability and will suffer less forgetting [112]. It must be noted that the performance of the first stage should be eliminated in order not to affect the forgetting evaluation. Hence, to make the methods share the same starting point, we utilize the same training strategy, data augmentation, and hyperparameters. We use basic data augmentation, e.g., random crop, horizontal flip, and color jitter for CIFAR100 and ImageNet.
It must be noted that Finetune, EWC, LwF, and L2P are exemplar-free methods, and we do not use any exemplar set for them. For other methods, we follow the benchmark setting to set the number of exemplars to 2,000 for CIFAR100 and ImageNet100 and 20,000 for ImageNet1000. These exemplars are equally sampled from each seen class via the herding [118] algorithm, as introduced in Definition 2.
4.2 Comparison on Small-Scale Datasets
In this section, we evaluate different methods on the small-scale dataset, i.e., CIFAR100. We compare these methods under both the TFS and TFH setting and design four data splits (two more splits are reported in supplementary). We report the incremental performance of different methods in Figure 6 and summarize the average and final performance, the number of parameters in Table II.
As we can infer from these figures, finetune shows the worst performance among all settings, verifying the fact that the model will suffer forgetting when sequentially learning new concepts. Regularizing the parameters, i.e., EWC, shows a negligible improvement over finetune. By contrast, LwF adds knowledge distillation loss to resist forgetting, which substantially improves the performance. When the exemplar set is available, directly replaying them during model updating can further enhance the performance by a substantial margin. Coil and iCaRL combine the exemplar repay and knowledge distillation and further obtain a performance boost than the vanilla replay. Model rectify methods, i.e., BiC and WA, rectify the bias in iCaRL and further improve the accuracy. However, recent methods based on dynamic networks (i.e., DER, FOSTER, MEMO, and DyTox) show competitive results with the help of multiple backbones. It indicates that saving more backbones can substantially help the model overcome forgetting. When it comes to pre-trained models, L2P obtains the best performance among all methods. However, other methods are trained from scratch, while L2P relies on the ViT pre-trained on ImageNet-21K, making it unfair to directly compare these two lines of methods.
@@ 4.3 Comparison on Large-Scale Datasets
In this section, we evaluate different methods on the large-scale dataset, i.e., ImageNet100 and ImageNet1000. We compare these methods under the TFS and TFH settings and design four data splits for ImageNet100 (two more splits are reported in supplementary) and two for ImageNet1000. We report the incremental performance (top-1 accuracy) of different methods in Figure 7 and Figure 8. We report the top-5 accuracy and summarize the average and final performance, the number of parameters in the supplementary. Note that GEM requires saving a large-scale matrix for solving the QP problem, which cannot be conducted with the ImageNet dataset. On the other hand, since the ViT in L2P is pre-trained on ImageNet-21K, incrementally training it on ImageNet is meaningless, and we do not report its results.
@@
Figure 11: Memory size of different comparison protocols. Dark bars denote the budget for saving exemplars, and red bars represent the budget for keeping the model in memory. Different methods should be aligned to the same budget for a fair comparison, as shown in (b).
4.4 Visualization of Confusion Matrix
Incremental models are biased toward the latest task, triggering the development of model rectify-based methods. In this section, we choose a typical phenomenon to show the model bias and how the CIL models rectify them, i.e., confusion matrix. We also visualize the weight norms of different methods in the supplementary.
Specifically, we show the confusion matrix of 10 typical methods in Figure 9 and report the full results in supplementary. The matrix is drawn under CIFAR100 Base0 Inc20 setting after the last incremental task. We can infer that finetune and EWC forget most knowledge from former classes, while LwF shows less forgetting with knowledge distillation. When equipping the model with exemplars, the diagonal elements become brighter, implying that forgetting is alleviated. Dynamic networks still show the most competitive performance in these visualizations, indicating that the former backbones help the model retrieve old features and resist catastrophic forgetting in the incremental learning process.
Table III: Average and last accuracy performance comparison on CIFAR100 with aligned memory cost. ‘#P’ represents the number of parameters (million). $\mathcal{E}$’ denotes the number of exemplars, and ‘MS’ denotes the memory size (MB). The full table is reported in the supplementary.
@@ 4.5 Memory-Aligned Comparison
As we can infer from former experimental evaluations, dynamic networks show the best performance among all methods. However, are these methods fairly compared? We argue that these dynamic networks implicitly introduce an extra memory budget, namely model buffer for keeping old models. The additional buffer results in an unfair comparison to those methods without storing models. Taking Table II for an example, the number of parameters in DER is ten times that of iCaRL in the CIFAR100 Base0 Inc10 setting. We visualize the memory cost of different methods in Figure 11(a) and find that dynamic networks obtain better performance at the expense of more memory budgets. It makes directly comparing different kinds of methods unfair.
@@
In this section, we follow [22] and aim for a fair comparison among different methods with different memory budgets. For those methods with different memory costs, we need to align the performance measure at the same memory scale for a fair comparison. For example, a ResNet32 model costs $463,504$ parameters (float), while a CIFAR image requires $3\times 32\times 32$ integer numbers (int). Hence, the budget for saving a backbone is equal to saving $463,504$ floats $\times 4$ bytes/float $\div(3\times 32\times 32)$ bytes/image $\approx 603$ instances for CIFAR. A fair comparison between different methods can be made by equipping the other methods with more exemplars. Since DER requires saving all the backbones from history, we align the memory cost of other methods to DER with extra exemplars. We visualize the memory budget of different methods under the current protocol in Figure 11(b), where the total budgets of different methods are aligned to the same scale.
@@
We report the results under the current protocol in Figure 10 and the detailed performance in Table III. Since Finetune, LwF, and EWC cannot be combined with the exemplar set, we do not report the results of these methods. Similarly, DyTox and L2P utilize different kinds of backbones, and we do not report their results. As we can infer from these results, the gap between dynamic networks and other methods is no longer large under fair comparison. For example, DER outperforms iCaRL by $9.07$ in terms of the final accuracy in the benchmark comparison of CIFAR100 Base0 Inc10, while the gap decreases to $1.74$ under the current protocol.
(a) CIFAR100 Base0 Inc10
(b) ImageNet100 Base50 Inc5
4.6 Memory-Agnostic Measure
To this end, we can set several ‘comparison budgets’ and align the memory cost of different methods to them. The budget list starts from a small value and incrementally enlarges, containing the requirement of different scale models. In this setting, we set the budget of the start point to a single backbone and gradually increase it to the budget cost of DER. For those algorithms without extra model storage, we equip more exemplars to meet the selected budget. However, when the selected budget is smaller than the required size, dynamic networks (i.e., DER and MEMO) cannot be deployed with the benchmark backbone. Therefore, we choose smaller backbones with fewer parameters for alignment. Hence, we can measure the performance of different methods at different memory scales and draw the performance-memory curve, as shown in Figure 12. The X-coordinate corresponds to the memory cost, and the Y-coordinate indicates the average performance $\bar{\mathcal{A}}$ (or last performance $\mathcal{A}_{B}$}). We suggest the area under the performance-memory curve (AUC) since the curve of each method indicates the dynamic ability with the change of model size. We calculate AUC-A and AUC-L, standing for the AUC under the average performance-memory and last performance-memory curves. The results are reported in Table IV.
As we can infer from these results, there are two main conclusions. Firstly, there exists an intersection between dynamic networks and other methods, where other methods perform better given a small memory size, while dynamic networks perform better given a large memory size. In other words, there is no free lunch in CIL, and different kinds of methods have their dominant domains. Secondly, the AUC measure provides a holistic way to measure the extendability of different methods given various memory budgets. We find that FOSTER and MEMO show competitive results regarding the AUC-A/L measure, implying their stronger extendability. It would be interesting to introduce it as the performance measure and design CIL methods for real-world applications. We report implementation details and the performance of each budget point in the supplementary.
4.7 Discussions about the Comparison
With the above experimental evaluations from three aspects, we have the following conclusions:
  • Equipping the model with exemplars is a simple and effective way to resist forgetting in CIL models.
  • Knowledge distillation performs better than parameter regularization in resisting forgetting with the same cost.
  • Model rectification can further boost the performance of other CIL models in a plug-and-play manner.
  • Pre-trained models can ease the burden of incremental learning and show very strong performance. However, since the features of pre-trained models are already available and do not need to be incrementally learned, comparing pre-trained models to other methods may be unfair.
  • Dynamic networks show the best performance in the evaluation at the cost of extra memory budgets. However, when changing the extra budgets into equal size of exemplars, the performance gap becomes smaller.
  • AUC-A and AUC-L provide a way to evaluate different CIL methods in a memory-agnostic manner, which can help select the method with extendability given any budget.
@@ 4.7 Discussions about the Comparison
Figure 13: Future directions of class-incremental learning. Class-incremental learning often emerges with other open-world problems, e.g., few-shot, long-tailed, multi-modal, weak supervision.

### v2

@@ 4 Experimental Evaluation
Specifically, Sections 4.2 and 4.3 answer the first question, and Section 4.4 answers the second question. We provide the holistic performance measures in Section 4.5 to answer the third question and summarize the results in Section 4.6. We report more results, measures, and visualizations in the supplementary material.
@@
(e) ImageNet100 Base0 Inc5
(f) ImageNet100 Base50 Inc10
(g) ImageNet1000 Base0 Inc100
(h) ImageNet1000 Base500 Inc100
Figure 6: Incremental performance of different methods on CIFAR100 (a-d), ImageNet100 (e-f), and ImageNet1000 (g-h). Legends are shown at the top of this figure, and we report the results of more settings in the supplementary.
@@ 4.1.1 Benchmark Datasets
iCaRL [82] firstly formulates the comparison protocol of class-incremental learning, which was widely followed and compared in other works. It suggests using CIFAR100 [24] and ImageNet100/1000 [207] for evaluation. CIFAR100 contains 100 classes with 60,000 images, in which 50,000 are training instances, and 10,000 are testing ones, with 100 images per class. Each image is represented by $\times32$ pixels. ImageNet1000 is a large-scale dataset with 1,000 classes, with about 1.28 million images for training and 50,000 for validation. ImageNet100 is the subset of ImageNet1000 containing 100 classes [82]. These classes are selected from the first 100 classes after a random shuffle. Some works [73, 72, 113, 27] utilize MNIST [208], CUB200 [209] and miniImageNet [207] for evaluation, while the aforementioned datasets are the most widely adopted in the current CIL community, and we choose them for model evaluation.
Table II: Average and last accuracy performance comparison on CIFAR100. ‘#P’ represents the number of parameters (million).
@@
Dataset Split: Following the protocol defined in [82], all classes are first shuffled by Numpy random seed $1993$. Afterward, there are two different ways to split the classes into incremental stages: $\bullet$ Train from scratch (TFS): splits all classes equally into each incremental stage. For example, if there are $B$ stages and $C$ classes in total, each incremental task contains C/B{C}/{B} classes for training.
$\bullet$ Train from half (TFH): splits half of the total classes as the first incremental task and equally assigns the rest classes into the following stages. Specifically, it assigns $C/2$ classes to the first task and C/2(B1)C/2(B-1) classes to the rest tasks.
Both of these settings are widely adopted in the current CIL community [89, 83]. Hence, we unify these settings as ‘Base-$m$, Inc-$n$’, where $m$ stands for the number of classes in the first stage, and $n$ stands for the number of classes in each incremental task. $m=0$ stands for the TFS protocol. We use the same training splits for every compared method for a fair comparison. The testing set is the same as the original one for holistic evaluation.
@@ 4.1.2 Evaluation Metrics
There are several metrics to evaluate the CIL model. We denote the Top-1 accuracy after the $b$-th task as $\mathcal{A}_{b}$, and higher $\mathcal{A}_{b}$ indicates a better prediction accuracy. Since the CIL model is continually updated, the accuracy often decays with more tasks incorporated. Hence, the accuracy after the last stage ($\mathcal{A}_{B}$) is a proper metric for measuring the overall accuracy among all classes.
@@
However, only comparing the final accuracy ignores the performance evolution along the learning trajectory. Hence, another metric denoted as ‘average accuracy’ considers the performance of all incremental stages: $\bar{\mathcal{A}}=\frac{1}{B}\sum_{b=1}^{B}\mathcal{A}_{b}$. A higher average accuracy denotes a stronger performance along the incremental stages. Apart from these measures, we also consider forgetting and intransigence [33] measures in the supplementary.
@@
4.1.3 Implementation Details
Selected methods: In the comparison, we aim to contain all kinds of methods in Table I. We systematically choose 17 methods, including: Replay [128], RMM [144] (data replay), GEM [53] (data regularization), EWC [73] (parameter regularization), AANets [66], FOSTER [16], MEMO [17], DER [15], DyTox [18], L2P [69] (dynamic networks), LwF [81], iCaRL [82], PODNET [94], Coil [88] (knowledge distillation), WA [112], BiC [83] (model rectify). We also report the baseline method ‘Finetune,’ which updates the model with Eq. 3.
The choice of these methods follows the development timeline of CIL, which is also the way we introduce these works. Additionally, it not only contains all seven aspects of CIL algorithms in our taxonomy but also includes early works (e.g., EWC, LwF, iCaRL) and recent state-of-the-art (DER, MEMO, L2P). The choice also gives consideration to CNN-based methods, ViT-based methods (DyTox), and even pre-trained ViT-based methods (L2P). Specifically, RMM is a specific technique to efficiently organize the memory budget, which can be orthogonally combined with other methods, and we combine it with FOSTER, denoted as FOSTER+RMM. Similarly, we combine AANets with LUCIR [89], denoted as LUCIR+AANets. L2P requires pre-trained ViT as the backbone model, while others are trained from scratch.
Training details: We implement the above methods with PyTorch [210] and PyCIL [211]. Specifically, we use the same network backbone for all CNN-based compared methods, i.e., ResNet32 [212] for CIFAR100 and ResNet18 for ImageNet. We use SGD with an initial learning rate of 0.1 and momentum of 0.9. The training epoch is set to 170 for all datasets with a batch size of 128. The learning rate suffers a decay of $0.1$ at 80 and 120 epochs. For ViT-based methods like DyTox and L2P, we follow the original implementation and use ConViT [213] for DyTox and pre-trained ViT-B/16 [19] for L2P. The optimization parameters of them are set according to the original paper since ViT has a different optimization preference to CNN. We follow the original paper to set the algorithm-specific parameters, e.g., splitting 10% exemplars from the exemplar set as validation for BiC, setting the temperature $\tau$ to 5 and using a 10 epochs warm-up for DER, using $\ell_{2}$ norm to normalize the fully-connected layers in WA. For EWC, the $\lambda$ parameter is done via a grid search among $\{1,10^{1},10^{2},10^{3},10^{4}\}$, and we find $10^{3}$ leads to its best performance.
Research has shown that a good starting point (i.e., the first stage accuracy $\mathcal{A}_{1}$) implies better transferability and will suffer less forgetting [110]. It must be noted that the performance gap of the first stage should be eliminated so as not to affect the forgetting evaluation. Hence, to make the methods share the same starting point, we utilize the same training strategy, data augmentation, and hyperparameters. We use basic data augmentation, e.g., random crop, horizontal flip, and color jitter for CIFAR100 and ImageNet.
It must be noted that Finetune, EWC, LwF, and L2P are exemplar-free methods, and we do not use any exemplar set for them. For other methods, we follow the benchmark setting to set the number of exemplars to 2,000 for CIFAR100 and ImageNet100 and 20,000 for ImageNet1000. These exemplars are equally sampled from each seen class via the herding [124] algorithm in Definition 2.
4.2 Comparison on Small-Scale Datasets
This section includes the comparison on the small-scale dataset, i.e., CIFAR100. We compare these methods under TFS and TFH settings with four data splits and report the incremental performance in Figure 6 (top). We summarize the average and final performance, the number of parameters in Table II.
As we can infer from these figures, finetune shows the worst performance among all settings, verifying the fact that the model will suffer forgetting when sequentially learning new concepts. Regularizing the parameters, i.e., EWC, shows a negligible improvement over finetune. By contrast, LwF adds knowledge distillation loss to resist forgetting, which substantially improves the performance. When the exemplar set is available, directly replaying them during model updating can further enhance the performance by a substantial margin. Coil and iCaRL combine the exemplar replay and knowledge distillation and further obtain a performance boost than the vanilla replay. Model rectify methods, i.e., BiC and WA, rectify the bias in iCaRL and further improve the accuracy. However, recent methods based on dynamic networks (i.e., DER, FOSTER, MEMO, and DyTox) show competitive results with the help of multiple backbones. It indicates that saving more backbones can substantially help the model overcome forgetting. When it comes to pre-trained models, L2P obtains the best performance among all methods. However, other methods are trained from scratch, while L2P relies on the ViT pre-trained on ImageNet-21K, making it unfair to directly compare these two lines of methods.
@@ 4.3 Comparison on Large-Scale Datasets
In this section, we evaluate different methods on the large-scale dataset, i.e., ImageNet100 and ImageNet1000. We compare these methods under the TFS and TFH settings and report the incremental performance (top-1 accuracy) of different methods in Figure 6 (bottom). We report the top-5 accuracy and summarize the average and final performance, the number of parameters in the supplementary. Since GEM requires saving a large-scale matrix for solving the QP problem, it cannot be conducted with the ImageNet dataset. On the other hand, since the ViT in L2P is pre-trained on ImageNet-21K, incrementally training it on ImageNet is meaningless, and we do not report its results.
@@
Figure 8: Memory size of different comparison protocols. Dark bars denote the budget for exemplars, and red bars represent the budget for keeping the model. Different methods should be aligned to the same budget for a fair comparison, as shown in (b).
Table III: Performance comparison on CIFAR100 with aligned memory cost. ‘#P’ represents the number of parameters (million). $\mathcal{E}$’ denotes the number of exemplars, and ‘MS’ denotes the memory size (MB).
@@ 4.4 Memory-Aligned Comparison
Former experimental evaluations indicate dynamic networks show the best performance among all methods. However, are these methods fairly compared? We argue that these dynamic networks implicitly introduce an extra memory budget, namely model buffer for keeping old models. The additional buffer results in an unfair comparison to those methods without storing models. Taking Table II for an example, the number of parameters in DER is ten times that of iCaRL in the CIFAR100 Base0 Inc10 setting. We visualize the memory cost of different methods in Figure 8(a) and find that dynamic networks obtain better performance at the expense of more memory budgets. It makes directly comparing different kinds of methods unfair.
@@
In this section, we follow [17] for a fair comparison among different methods with different memory budgets. For those methods with different memory costs, we need to align the performance measure at the same memory scale for a fair comparison. For example, a ResNet32 model costs $463,504$ parameters (float), while a CIFAR image requires $3\times 32\times 32$ integer numbers (int). Hence, the budget for saving a backbone is equal to saving $463,504$ floats $\times 4$ bytes/float $\div(3\times 32\times 32)$ bytes/image $\approx 603$ instances for CIFAR. A fair comparison between different methods can be made by equipping the other methods with more exemplars. Since DER requires saving all the backbones from history, we align the memory cost of other methods to DER with extra exemplars. We visualize the memory budget of different methods under the current protocol in Figure 8(b), where the total budgets of different methods are aligned to the same scale.
@@
We report the results under the current protocol in Figure 7 and the detailed performance in Table III. Since Finetune, LwF, and EWC cannot be combined with the exemplar set, we do not report the results of these methods. Similarly, DyTox and L2P utilize different kinds of backbones, and we do not report their results. As we can infer from these results, the gap between dynamic networks and other methods is no longer large under fair comparison. For example, DER outperforms iCaRL by 9.079.07% in terms of the final accuracy in the benchmark comparison of CIFAR100 Base0 Inc10, while the gap decreases to 1.741.74% under the current protocol.
@@
(a) CIFAR100 Base0 Inc10
(b) ImageNet100 Base50 Inc5
4.5 Memory-Agnostic Measure
To this end, we can set several ‘comparison budgets’ and align the memory cost of different methods to them. The budget list starts from a small value and incrementally enlarges, containing the requirement of different scale models. In this setting, we set the budget of the start point to a single backbone and gradually increase it to the budget cost of DER. For those algorithms without extra model storage, we equip more exemplars to meet the selected budget. However, when the selected budget is smaller than the required size, dynamic networks (i.e., DER and MEMO) cannot be deployed with the benchmark backbone. Therefore, we choose smaller backbones with fewer parameters for alignment. Hence, we can measure the performance of different methods at different memory scales and draw the performance-memory curve, as shown in Figure 9. The X-coordinate corresponds to the memory cost, and the Y-coordinate indicates the average performance $\bar{\mathcal{A}}$ (or last performance $\mathcal{A}_{B}$}). We suggest the area under the performance-memory curve (AUC) since the curve of each method indicates the dynamic ability with the change of model size. We calculate AUC-A and AUC-L, standing for the AUC under the average performance-memory and last performance-memory curves.
As we can infer from Table IV, there are two main conclusions. Firstly, there exists an intersection between dynamic networks and other methods, where other methods perform better given a small memory size. In contrast, dynamic networks perform better given a large memory size. In other words, there is no free lunch in CIL, and different kinds of methods have their dominant domains. Secondly, the AUC measure provides a holistic way to measure the extendability of different methods given various memory budgets. FOSTER and MEMO show competitive results regarding the AUC-A/L measure, implying their stronger extendability. It would be interesting to introduce it as the performance measure and design CIL methods for real-world applications. We report implementation details and the performance of each budget point in the supplementary.
4.6 Discussions about the Comparison
With the above experimental evaluations from three aspects, we have the following conclusions: 1) Equipping the model with exemplars is a simple and effective way to resist forgetting in CIL models. 2) Knowledge distillation performs better than parameter regularization in resisting forgetting with the same cost. 3) Model rectification can further boost the performance of other CIL models in a plug-and-play manner. 4) Pre-trained models can ease the burden of incremental learning and show very strong performance. However, since the features of pre-trained models are already available and do not need to be incrementally learned, comparing pre-trained models to other methods may be unfair. 5) Dynamic networks show the best performance in the evaluation at the cost of extra memory budgets. However, when changing the extra budgets into equal size of exemplars, the performance gap becomes smaller. 6) AUC-A and AUC-L provide a way to evaluate different CIL methods in a memory-agnostic manner, which can help select the method with extendability given any budget.